OpenStudy (anonymous):

Bounded by y=(3x^2+x)/((x^2+1)(x+1)), y=0, x=1

OpenStudy (tkhunny):

What is to be done?

OpenStudy (tkhunny):

You may wish to try Partial Fractions! \(\dfrac{3x^{2} + x}{(x^{2}+1)\cdot (x+1)} = \dfrac{1}{x+1} + \dfrac{2x}{x^{2}+1} + \dfrac{-1}{x^{2}+1}\) Do you know how to find that expression? There are three separate integration techniques to exercise, there.